# How Uniform Random Weights Induce Non-uniform Bias: Typical Interpolating Neural Networks Generalize with Narrow Teachers

### v1

@@ Abstract
Background. A main theoretical puzzle is why over-parameterized Neural Networks (NNs) generalize well when trained to zero loss (i.e., so they interpolate the data). Usually, the NN is trained with Stochastic Gradient Descent (SGD) or one of its variants. However, recent empirical work examined the generalization of a random NN that interpolates the data: the NN was sampled from a seemingly uniform prior over the parameters, conditioned on that the NN perfectly classifying the training set. Interestingly, such a NN sample typically generalized as well as SGD-trained NNs.
Contributions. We prove that such a random NN interpolator typically generalizes well if there exists an underlying narrow “teacher NN" that agrees with the labels. Specifically, we show that such a ‘flat’ prior over the NN parametrization induces a rich prior over the NN functions, due to the redundancy in the NN structure. In particular, this creates a bias towards simpler functions, which require less relevant parameters to represent — enabling learning with a sample complexity approximately proportional to the complexity of the teacher (roughly, the number of non-redundant parameters), rather than the student’s.
@@ 1 Introduction
A central theoretical question in deep learning is why Neural Networks (NNs) generalize, despite being over-parameterized, and even when perfectly fitted to the data (Zhang et al., 2017). One of the leading explanations for this phenomenon is that NNs have an “implicit bias" toward generalizing solutions (e.g., Gunasekar et al. (2017); Soudry et al. (2018); Arora et al. (2019); Lyu and Li (2020); Chizat and Bach (2020); Vardi (2023)). This bias stems from underlying interactions between the model and the training method — including the type of optimization step, the initialization, the parametrization, and the loss function.
@@
These results may suggest that such a uniform sampling of the NN parameters induces simple and generalizing NN functions. In this paper, we prove this is indeed the case, and aim to uncover the mechanism behind this phenomenon. In short, we prove typical NN interpolators sampled this way (“students") generalize well, given there exists a “narrow” NN teacher that generates the labels. In more detail,
  • In Section 3 we prove that a typical NN sampled from the posterior over interpolators generalizes well (i.e., has a small test error with high probability) with
    #$\#\textrm{samples}=O\left(-\log\tilde{p}\right)\,,$
    where $\tilde{p}$ is the probability that a random NN (sampled from the ‘uniform’ prior) is equivalent to the teacher function. [^2] Proof idea: the number of hypotheses sampled until a successful interpolation is $\left|\mathcal{H}\right|\lessapprox 1/\tilde{p}$. Plugging this into the standard sample complexity of a finite hypotheses class $O\left(\log\left|\mathcal{H}\right|\right)$, we obtain the result. The actual proof is slightly more complicated since $\left|\mathcal{H}\right|$ here weakly depends on the training set. Thus, to obtain generalization guarantees for NNs, we proceed to upper bound $\left(-\log\tilde{p}\right)$.
@@
    • The proofs in both cases are simple [^3] Proof idea, for a two-layer FC net without biases: The NN function is identical to the teacher NN function, if we set $d_{1}^{\star}$ hidden neurons with the same ingoing and outgoing weights as in the teacher; and for the other $d_{1}-d_{1}^{\star}$ hidden neurons we set the outgoing weights (for the zeroed neurons, the input weights do not matter). The probability of this is $\tilde{p}=Q^{-d_{0}d_{1}^{\star}-d_{1}}$, which satisfies (1)., and can be extended for more general architectures.
  • Combining these relatively easy-to-prove results, we get a surprisingly novel result: typical NN interpolators have sample complexity approximately proportional to the number of teacher parameters times the number of quantization bits, with only a weak dependence on the student width. Thus, the student generalizes well if there exists a teacher that is sufficiently narrow and under-parameterized in comparison to the sample number.
  • Finally, in Section 6 we examine a two-layer neural network with continuous weights and derive similar results, except a margin factor replaces $Q$ in the bound.
@@ 2 Preliminaries
Hypothesis parametrization. We discuss parameterized predictors $\bm{\theta}\mapsto h_{\bm{\theta}}$, where $\bm{\theta}\in\mathbb{R}^{M}$. Distributions over $\bm{\theta}$ therefore induce distributions over hypotheses via
$\mathcal{P}\left(h\right)\triangleq\mathbb{P}_{\bm{\theta}}\left(\bm{\theta}\mid h_{\bm{\theta}}=h\right)~{}.$
@@ 3 Generalization Bounds for Random Interpolating Hypotheses
$\mathcal{A}_{\mathcal{P}}(\mathcal{S})\sim\mathcal{P}_{S},$ (3)
@@
We will be particularly interested in the case in which $\mathcal{P}(h)$ is defined through a ‘uniform’ (or otherwise fairly ‘flat’ or benign) prior on the parameters $\bm{\theta}$ in some parametrization $h_{\bm{\theta}}$. But in this section, we analyze posterior sampling, or equivalently Guess and Check, directly through the induced distribution $\mathcal{P}$ over predictors. In particular, we analyze generalization performance in terms of the probability that a random hypothesis $h\sim\mathcal{P}$ is equivalent to the teacher model. Formally:
@@
We say that a predictor $h$ is teacher-equivalent (TE) w.r.t. a data distribution $\mathcal{D}$, and denote $h\equiv h^{\star}$, if $\mathbb{P}_{\mathbf{x}\sim\mathcal{D}}\left(h\left(\mathbf{x}\right)=h^{\star}\left(\mathbf{x}\right)\right)=1\,.$ and denote the probability of a random hypothesis to be TE by
@@
As we show in the next result, $\tilde{p}$ plays an important role in G&C generalization. Specifically, in Appendix A.2 we derive the following generalization bound.
Theorem 3.2 (G&C (i.e. Posterior Sampling) Generalization).
Let $\epsilon\in\left(0,1\right)$ and $\delta\in\left(0,\frac{1}{5}\right)$, and assume that $\tilde{p}<\frac{1}{2}$. For any $N$ larger than
$\frac{-\log\left({\tilde{p}}\right)+3\log\left(\frac{2}{\delta}\right)}{\epsilon}\,,$
@@
$\mathbb{P}_{S\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)<\epsilon\right)\geq 1-\delta\,.$
We observe that the sample complexity required to ensure $(\epsilon,\delta)$-PAC generalization depends on $\left(-\log\left(\tilde{p}\right)\right)$. Thus, we define the effective sample complexity as
@@
Moreover, we notice the above theorem implies (from Markov inequality, see Appendix A.3):
@@
For $\epsilon,\delta$ as above, and any $N$ larger than
$\frac{-\log\left({\tilde{p}}\right)+6\log\left(\frac{2}{\delta}\right)}{\epsilon}\,,$
@@
$\mathbb{P}_{\mathcal{S}}\left(\mathbb{P}_{h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\geq\epsilon\right)<\delta\right)\geq 1-\delta\,.$
@@
Corollary 3.3 implications. Corollary 3.3 examines, for a single sample of the data $\mathcal{S}$, the relative volume of ‘bad’ interpolators out of all interpolators — i.e. the probability to sample an interpolator for which $\mathcal{L_{D}}\left(h\right)\geq\epsilon$, given the data $\mathcal{S}$. It states that this relative volume is small ($\delta$) with high probability ($1-\delta$) over the sampling of the data. And $\delta$ can be quite small, since for any $\epsilon$, we have that $\delta$ decays exponentially fast in $N$
$\delta=2\exp\left(-\frac{\epsilon N+\log\left(\tilde{p}\right)}{6}\right).$
@@
Proof idea of Theorem 3.2 and relationship to PAC-Bayes. In Appendix A.2, we prove Theorem 3.2 by noting that the expected number of hypotheses we will consider is $1/\tilde{p}$, and so we are essentially selecting an interpolating hypothesis from the effective hypothesis class =$\mathcal{H}=\{h_{1},\ldots,h_{\tau}\}$ with $\tau\approx 1/\tilde{p}$, where the hypotheses in this class are chosen before seeing the training set. The sample complexity is thus $\log\left|\mathcal{H}\right|=\log\tau\approx-\log\tilde{p}$. The only complication is that the stopping time $\tau$ of G&C is random and depends on $\mathcal{S}$. But it is enough to bound $\tau$ very crudely (which we do with high probability), as a multiplicative factor to $\tau$ results only in an additive logarithmic factor.
@@
The analysis here is similar to PAC-Bayes analysis (McAllester, 1999), which also studies the behavior of a posterior over hypotheses, except that in typical PAC-Bayes analysis the bound is over the expected population error $\mathbb{E}_{h\sim\mathcal{P}_{\mathcal{S}}}\left[\mathcal{L}_{\mathcal{D}}(h)\right]$ for a sample from the posterior. In fact, noting that $\mathrm{KL}(\mathcal{P}_{\mathcal{S}}\|\mathcal{P})=-\log\mathbb{P}_{h\sim\mathcal{P}}(\mathcal{L}_{\mathcal{S}}(h)\!=\!0)\geq-\log\tilde{p}$, a standard PAC-Bayes bound (Langford and Seeger, 2002; McAllester, 2003) will yield that with the same sample complexity as in Theorem 3.2,
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\;\mathbb{E}_{h\sim\mathcal{P}_{\mathcal{S}}}\left[\mathcal{L}_{\mathcal{D}}(h)\right]<\epsilon\;\right)\geq 1-\delta.$ (4)
@@
The difference is that Theorem 3.2 holds with high probability for a single posterior sample, instead of just in the expectation over the posterior. That is, Theorem 3.2 establishes that not only are random interpolators good on average, but only a small fraction of them are bad.
@@
Using Markov inequality one can derive from (4) a high probability bound for a for a single draw from the posterior (as in Theorem 3.2), but that bound is less tight then Theorem 3.2 (see Appendix A.4). Also, a variant of the PAC-Bayes theorem that holds with high probability for a single draw from the posterior has also been presented by Alquier (2023, Theorem 2.7). As is, Alquier’s result yields a much looser bound, since it is more generic (it applies to any posterior, not just conditioning on interpolation). Theorem 3.2 can be viewed as a tighter specialization to interpolators.
A sample complexity of $\left(-\log\tilde{p}\right)$ should not be surprising, and can also be obtained by a Minimum Description Length (MDL)/Occam Razor learning rule $\textsc{MDL}_{\mathcal{P}}(\mathcal{S})=\arg\max_{\mathcal{L}_{\mathcal{S}}(h)=0}\mathcal{P}(h)$ (Blumer et al., 1987) (see also Section 7.3 in Shalev-Shwartz and Ben-David (2014)). But importantly, this requires using the induced probability mass function over hypotheses, and not the distribution over parametrizations—see discussion in the next Section.
4 Redundancy and Description Length
Before we state our main results for NNs in Section 5, in this section, we pause to give a broader perspective on the effect of the parametrization on $\tilde{p}=\mathbb{P}_{h\sim\mathcal{P}}\left(h\equiv h^{\star}\right)$. As we saw, the parametrization controls the generalization behavior of G&C and posterior sampling, and how a seemingly uniform prior over parameters can induce a rich prior over hypothesis via redundancy. We first consider a conceptually very direct parametrization.
How parameter redundancy works: a minimal example. Consider a predictor $h_{\sigma}$ specified by a description language (i.e. parametrization) using bit strings $\sigma$, where descriptions end with the string “END” (encoded in bits), and this string never appears elsewhere in the description [^4] This can be a Turing complete programming language, or perhaps better to think of simpler descriptions such as boolean formulas or strings encoding decision trees.. We will consider students that use fixed-length descriptions, i.e., $\sigma\in\{0,1\}^{C}$, where we only consider the description up to the first “END”, i.e. $h_{\sigma}=h_{\tilde{\sigma}}$ where $\tilde{\sigma}$ is a prefix of $\sigma$ ending with the first “END”. [^5] If $\sigma$ does not contain “END”, or $\tilde{\sigma}$ is not a valid description, we can set $h_{\sigma}$ to the constant 0 predictor. Let us now consider a uniform prior over this parametrization, i.e. $\mathcal{P}(\sigma)=2^{-C}$ for all $\sigma\in\{0,1\}^{C}$. Although this prior is uniform over parametrizations, it is easy to see that it induces a highly non-uniform prior over predictors: for a predictor $h_{\tau}$ with a short description $\tau\in\{0,1\}^{{\left\lvert{\tau}\right\rvert}}$, ${\left\lvert{\tau}\right\rvert}\ll C$, it’s $C$-bit description is highly redundant, as the final $C-{\left\lvert{\tau}\right\rvert}$ bits can be set arbitrarily, as long as the first bits are $\tau$, and so $\mathcal{P}(h_{\tau})\geq 2^{-{\left\lvert{\tau}\right\rvert}}$. For a teacher with a short description $h^{\star}=h_{\tau^{\star}}$, $\tau^{\star}\in\{0,1\}^{C^{\star}}$, we thus have $-\log\tilde{p}\leq C^{\star}$. The sample complexity of learning a small teacher thus depends only on the complexity of the teacher and not the complexity of the (possibly much larger) student.
Why would MDL fail here? Returning to the MDL/Occam principle mentioned at the end of Section 3, it is important to note that applying this principle to the distribution over parameters would not work here. Consider the MDL/Occam/MAP rule which selects the most likely interpolating parameters $\arg\max_{\mathcal{L}_{\mathcal{S}}(h_{\sigma})}\mathcal{P}(\sigma)$. This rule would be useless here, since for all $\sigma$ in our parameter space $\{0,1\}^{C}$ have the same prior probability, and its associated sample complexity would be $O(-\log\mathcal{P}(\sigma^{\star}))=C$ (not $\mathcal{P}(h_{\sigma^{\star}})$ !), where $\sigma^{\star}$ is a description of $h^{\star}$ in our parameter space (e.g. $\sigma^{\star}=\tau^{\star}+$‘0’$\cdot(C-C^{\star})$). The important distinction is that Occam considers the probability mass function over parameters (roughly, the density under a specific parametrization or base measure). On the other hand, even if we sample $h_{\sigma}$ by sampling parameters $\sigma$, posterior sampling (i.e. G&C) can be thought of directly in terms of the distribution over hypotheses.
When would redundancy fail? To see how changing the parametrization can change the induced distribution over hypothesis, consider instead a non-redundant parametrization, where we take $\sigma$ to be uniform over valid descriptions, of length at most $C$, ending with “END” (i.e. strings ending with “END” and that do not otherwise contain “END”). The induced prior over hypothesis is now uniform [^6] We might still have further redundancies in that multiple valid descriptions can describe the same function, again introducing non-uniformity. Consider here a non-redundant description language, e.g. a read-once branching program. and even using posterior sampling/G&C would have sample complexity determined by the complexity of the student, rather than the teacher.
When would MDL succeed? Finally, we note that we could of course learn a short teacher with an Occam rule that is explicitly biased towards short descriptions, e.g. using the Kraft prior [^7] We again absorb any remaining probability in the constant zero predictor. Here we are thinking of an unbounded student and the parameter space being any string ending with the first occurrence of “END”. We can also of course bound the size of the student. $\mathcal{P}(\sigma)=2^{-{\left\lvert{\sigma}\right\rvert}}$. Using such a prior, the Occam rule also enjoys a sample complexity of $O(C^{\star})$ that depends only on the length of the teacher. But here the prior over parameters is non-uniform and explicitly biased, and our interest in this paper is in how seemingly uniform priors over parameters can induce non-uniform priors and generalization over predictors due to the choice of parametrization.

### v2

@@ Abstract
Background. A main theoretical puzzle is why over-parameterized Neural Networks (NNs) generalize well when trained to zero loss (i.e., so they interpolate the data). Usually, the NN is trained with Stochastic Gradient Descent (SGD) or one of its variants. However, recent empirical work examined the generalization of a random NN that interpolates the data: the NN was sampled from a seemingly uniform prior over the parameters, conditioned on that the NN perfectly classifies the training set. Interestingly, such a NN sample typically generalized as well as SGD-trained NNs. Contributions. We prove that such a random NN interpolator typically generalizes well if there exists an underlying narrow “teacher NN” that agrees with the labels. Specifically, we show that such a ‘flat’ prior over the NN parameterization induces a rich prior over the NN functions, due to the redundancy in the NN structure. In particular, this creates a bias towards simpler functions, which require less relevant parameters to represent — enabling learning with a sample complexity approximately proportional to the complexity of the teacher (roughly, the number of non-redundant parameters), rather than the student’s.
Machine Learning, ICML
@@ 1 Introduction
A central theoretical question in deep learning is why Neural Networks (NNs) generalize, despite being over-parameterized, and even when perfectly fitted to the data (Zhang et al., 2017). One of the leading explanations for this phenomenon is that NNs have an “implicit bias” toward generalizing solutions (e.g., Gunasekar et al. (2017); Soudry et al. (2018); Arora et al. (2019); Lyu & Li (2020); Chizat & Bach (2020); Vardi (2023)). This bias stems from underlying interactions between the model and the training method — including the type of optimization step, the initialization, the parameterization, and the loss function.
@@
These results may suggest that such a uniform sampling of the NN parameters induces simple and generalizing NN functions. In this paper, we prove this is indeed the case, and aim to uncover the mechanism behind this phenomenon. In short, we prove typical NN interpolators sampled this way (“students”) generalize well, given there exists a “narrow” NN teacher that generates the labels. Next, we explain these results in more detail.
Contributions. In Section 3 we prove that a typical NN sampled from the posterior over interpolators generalizes well (i.e., has a small test error with high probability) with
#$\#\textrm{samples}=O\left(-\log\tilde{p}\right)\,,$ (1)
where $\tilde{p}$ is the probability that a random NN (sampled from the ‘uniform’ prior) is equivalent to the teacher function. [^2] The proof idea is simple: the number of hypotheses sampled until a successful interpolation is $\left|\mathcal{H}\right|\lessapprox 1/\tilde{p}$. Plugging this into the standard sample complexity of a finite hypotheses class $O\left(\log\left|\mathcal{H}\right|\right)$, we obtain the result. The actual proof is slightly more complicated since $\left|\mathcal{H}\right|$ here weakly depends on the training set. Thus, to obtain generalization guarantees for NNs, we proceed to upper bound $\left(-\log\tilde{p}\right)$.
@@
  • The proofs in both cases are simple [^3] Proof idea for two-layer FC nets without biases: The NN function is identical to the teacher NN function, if we set $d_{1}^{\star}$ hidden neurons with the same ingoing and outgoing weights as in the teacher; and for the other $d_{1}-d_{1}^{\star}$ hidden neurons we set the outgoing weights (for the zeroed neurons, the input weights do not matter). This event probability is $\tilde{p}=Q^{-d_{0}d_{1}^{\star}-d_{1}}$, satisfying (2)., and can be extended for more general architectures.
Lastly, in Section 5 we examine a two-layer neural network with continuous weights and derive similar results, except a margin assumption replaces the quantization assumption, and a margin factor replaces $Q$ in the bound.
Implications. Combining these relatively easy-to-prove results ((1) and (2)), we get a surprisingly novel result: typical NN interpolators have sample complexity approximately proportional to the number of teacher parameters times the number of quantization bits, with only a weak dependence on the student width. Thus, the student generalizes well if there exists a teacher that is sufficiently narrow and under-parameterized in comparison to the sample number. As a corollary, we show that with high probability over the training set, the volume of interpolators with high generalization error is exponentially small in the size of the training set.
In Section 7 we discuss our assumptions (teacher narrowness and weight quantization), how our results can be straightforwardly extended beyond interpolators (to functions with a non-zero training error), whether posterior sampling biases us towards sparse representations, the effect of parameterization via the minimum description length framework, and the relation of our results to SGD.
@@ 2 Preliminaries
Hypothesis parameterization. We discuss parameterized predictors $\boldsymbol{\theta}\mapsto h_{\boldsymbol{\theta}}$, where $\boldsymbol{\theta}\in\mathbb{R}^{M}$. Distributions over $\boldsymbol{\theta}$ therefore induce distributions over hypotheses via
$\mathcal{P}\left(h\right)\triangleq\mathbb{P}_{\boldsymbol{\theta}}\left(h_{\boldsymbol{\theta}}=h\right)~{}.$
That is, $\mathcal{P}(h)$ is the probability mass function of sampling parameters $\boldsymbol{\theta}$ mapping to $h$ when the distribution is discrete, or (with a slight abuse of notation) their density when the distribution is continuous.
@@ 3 Generalization Bounds for Random Interpolating Hypotheses
$\mathcal{A}_{\mathcal{P}}(\mathcal{S})\sim\mathcal{P}_{\mathcal{S}}\,,$ (4)
@@
We will be particularly interested in the case in which $\mathcal{P}(h)$ is defined through a ‘uniform’ (or otherwise fairly ‘flat’ or benign) prior on the parameters $\boldsymbol{\theta}$ in some parameterization $h_{\boldsymbol{\theta}}$. But in this section, we analyze posterior sampling, or equivalently Guess and Check, directly through the induced distribution $\mathcal{P}$ over predictors. In particular, we analyze generalization performance in terms of the probability that a random hypothesis $h\sim\mathcal{P}$ is equivalent to the teacher model. This is formalized in the following definition.
@@
We say that a predictor $h$ is teacher-equivalent (TE) w.r.t. a data distribution $\mathcal{D}$, and denote $h\equiv h^{\star}$, if $\mathbb{P}_{\mathbf{x}\sim\mathcal{D}}\left(h\left(\mathbf{x}\right)=h^{\star}\left(\mathbf{x}\right)\right)=1\,,$ and denote the probability of a random hypothesis to be TE by
@@
As we show in the next result, $\tilde{p}$ plays an important role in G&C generalization. Specifically, in Appendix B.2 we derive the following generalization bound.
Lemma 3.2 (G&C (i.e. Posterior Sampling) Generalization).
Let $\varepsilon\in\left(0,1\right)$ and $\delta\in\left(0,\frac{1}{5}\right)$, and assume that $\tilde{p}<\frac{1}{2}$. For any $N$ larger than
$\frac{-\log\left({\tilde{p}}\right)+3\log\left(\frac{2}{\delta}\right)}{\varepsilon}\,,$
@@
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)<\varepsilon\right)\geq 1-\delta\,.$
We observe that the sample complexity required to ensure $(\varepsilon,\delta)$-PAC generalization depends on $\left(-\log\left(\tilde{p}\right)\right)$. Thus, we define the effective sample complexity as
@@
Moreover, using Markov’s inequality, the above lemma implies (see Appendix B.3) the following corollary.
@@
For $\varepsilon,\delta$ as above, and any $N$ larger than
$\frac{-\log\left({\tilde{p}}\right)+6\log\left(\frac{2}{\delta}\right)}{\varepsilon}\,,$
@@
$\mathbb{P}_{\mathcal{S}}\left(\mathbb{P}_{h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\geq\varepsilon\right)<\delta\right)\geq 1-\delta\,.$
@@
Corollary 3.3 implications. Corollary 3.3 examines, for a single sample of the data $\mathcal{S}$, the relative volume of ‘bad’ interpolators out of all interpolators — i.e. the probability to sample an interpolator for which $\mathcal{L_{D}}\left(h\right)\geq\varepsilon$, given the data $\mathcal{S}$. It states that this relative volume is small ($\delta$) with high probability ($1-\delta$) over the sampling of the data. And $\delta$ can be quite small, since for any $\varepsilon$, we have that $\delta$ decays exponentially fast in $N$
$\delta=2\exp\left(-\frac{\varepsilon N+\log\left(\tilde{p}\right)}{6}\right).$
@@
Proof idea of Lemma 3.2 and relationship to PAC-Bayes. In Appendix B.2, we prove Lemma 3.2 by noting that the expected number of hypotheses we will consider is $1/\tilde{p}$, and so we are essentially selecting an interpolating hypothesis from the effective hypothesis class =$\mathcal{H}=\{h_{1},\ldots,h_{\tau}\}$ with $\tau\approx 1/\tilde{p}$, where the hypotheses in this class are chosen before seeing the training set. The sample complexity is thus $\log\left|\mathcal{H}\right|=\log\tau\approx-\log\tilde{p}$. The only complication is that the stopping time $\tau$ of G&C is random and depends on $\mathcal{S}$. But it is enough to bound $\tau$ very crudely (which we do with high probability), as a multiplicative factor to $\tau$ results only in an additive logarithmic factor.
Remark 3.4.
The same method can be straightforwardly used to extend these results to the case where the teacher NN is not a perfect interpolator, and the G&C algorithm is modified to stop when the training error is bellow some threshold instead of 0 (see Appendix B.4).
@@
The analysis here is similar to PAC-Bayes analysis (McAllester, 1999), which also studies the behavior of a posterior over hypotheses, except that in typical PAC-Bayes analysis the bound is over the expected population error $\mathbb{E}_{h\sim\mathcal{P}_{\mathcal{S}}}\left[\mathcal{L}_{\mathcal{D}}(h)\right]$ for a sample from the posterior. In fact, noting that $\mathrm{KL}(\mathcal{P}_{\mathcal{S}}\|\mathcal{P})=-\log\mathbb{P}_{h\sim\mathcal{P}}(\mathcal{L}_{\mathcal{S}}(h)\!=\!0)\geq-\log\tilde{p}$, a standard PAC-Bayes bound (Langford & Seeger, 2002; McAllester, 2003) will yield that with the same sample complexity as in Lemma 3.2,
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\;\mathbb{E}_{h\sim\mathcal{P}_{\mathcal{S}}}\left[\mathcal{L}_{\mathcal{D}}(h)\right]<\varepsilon\;\right)\geq 1-\delta.$ (5)
@@
The difference is that Lemma 3.2 holds with high probability for a single posterior sample, instead of just in the expectation over the posterior. That is, Lemma 3.2 establishes that not only are random interpolators good on average, but only a small fraction of them are bad.
@@
Using Markov’s inequality one can derive from (5) a high probability bound for a for a single draw from the posterior (as in Lemma 3.2), but that bound is less tight then Lemma 3.2 (see Appendix B.5). Also, a variant of the PAC-Bayes theorem that holds with high probability for a single draw from the posterior has also been presented by Alquier (2023, Theorem 2.7). As is, Alquier’s result yields a much looser bound, since it is more generic (it applies to any posterior, not just conditioning on interpolation). Thus, Lemma 3.2 can be viewed as a tighter specialization to interpolators.
A sample complexity of $(-\log\tilde{p})$ should not be surprising, and can also be obtained by an Occam Razor / Minimum Description Length learning rule $\mathrm{MDL}_{\mathcal{P}}(\mathcal{S})=\arg\max_{\mathcal{L}_{\mathcal{S}}(h)=0}\mathcal{P}(h)$ (Blumer et al. (1987), and see also Section 7.3 in Shalev-Shwartz & Ben-David (2014)). Here, we discussed how the same sample complexity is obtained by a single draw from the posterior, as in G&C. More interesting is how, starting from a uniform prior $\mathcal{P}_{\boldsymbol{\theta}}(\boldsymbol{\theta})$ over parameters, we end up with an informative induced prior $\mathcal{P}(h)$ over hypotheses, which has high $\tilde{p}$ and thus low sample complexity. The key here is redundancy in the parameterization. In Appendix A we review the general principle of how non-uniform redundancy in the parameterization can induce non-uniform informative priors $\mathcal{P}(h)$ and thus low sample complexity. In the next sections we see how this plays specifically for NNs, analyzing $\tilde{p}$ under the prior induced by a uniform choice of NN parameters.